Applied Overhead and Unit Overhead Cost: Plantwide Rates

Seco, Inc., produces two types of clothes dryers: deluxe and regular. Seco uses a plantwide rate based on direct labor hours to assign its overhead costs. The company has the following estimated and actual data for the coming year:

Estimated overhead $1,664,000
Expected activity 52,000
Actual activity (direct labor hours):
Deluxe dryer 12,000
  Regular dryer 40,000
Units produced:
Deluxe dryer 24,000
  Regular dryer 200,000

Required:

1. Calculate the predetermined plantwide overhead rate, using direct labor hours.
$ per hour

Calculate the applied overhead for each product, using direct labor hours.

Applied overhead
Deluxe $
Regular $

2. Calculate the overhead cost per unit for each product. If required, round your answers to the nearest cent.

Overhead Cost
Deluxe $ per unit
Regular $ per unit

3. What if the deluxe product used 24,000 hours (to produce 24,000 units) instead of 12,000 hours (total expected hours remain the same)? Calculate the effect on the profitability of this product line if all 24,000 units are sold.

Profits would - Select your answer -increase decrease by $
Answer #1

1.

Estimated overhead = $1,664,000

Estimated direct labor hours = 52,000

Predetermined plant wide overhead rate = Estimated overhead/Estimated direct labor hours

= 1,664,000/52,000

= $32 per direct labor hour

Overhead applied to Deluxe = Direct labor hours used by Deluxe x Predetermined plant wide overhead rate

= 12,000 x 32

= $384,000

Overhead applied to Regular = Direct labor hours used by Deluxe x Predetermined plant wide overhead rate

= 40,000 x 32

= $1,280,000

2.

Overhead cost per unit of Deluxe = Overhead applied to Deluxe/Number of units of Deluxe

= 384,000/24,000

= $16

Overhead cost per unit of Regular = Overhead applied to Regular/Number of units of Regular

= 1,280,000/200,000

= $6.4

3.

Overhead applied to Deluxe = Direct labor hours used by Deluxe x Predetermined plant wide overhead rate

= 24,000 x 32

= $768,000

Overhead cost per unit of Deluxe = Overhead applied to Deluxe/Number of units of Deluxe

= 768,000/24,000

= $32

Overhead cost per unit of Deluxe would increase by $16 (32 - 16)

Profits would decrease by = 24,000 x 16

= $384,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.